SÉMINAIRE DE GÉOMÉTRIE

Vendredi, 14 Janvier 2000 10:30 – 11:30 Salle 2 , Laboratoire MPB

Résumé :

A closed Riemann surface is called maximal if the length of its shortest closed geodesic is a local maximum in Teichmueller space. A triangle surface is a surface whose automorphism group is a nontrivial finite quotient of a triangle group. We give a classification of all maximal triangle surfaces of a particular simple type.